The number of ladies is 80 greater than the number of men. If there are 1310 people all together, how many gentleman are there?
Answer by Maths68(1474)

Ladies = l
Men = m
The number of ladies is 80 greater than the number of men
l=m+80
there are 1310 people all together
l+m=1310
Put the value of l in above equation
(m+80)+m=1310
2m+80=1310
2m=1310-80
2m=1230
2m/2=1230/2
m=615
Given
l=m+80
l=615+80
l=695
Ladies = l = 695
Men = m = 615